Hi! I was given a problem such as this: "Write a polynomial function f of least degree that has rational coefficients, a leading coefficient of 1, and -7, -4 as the given zeros."
How am I supposed to do it?

Answer by jim_thompson5910(35256)

x = -7 or x = -4

x+7 = 0 or x+4 = 0

(x+7)(x+4) = 0

x^2 + 11x + 28 = 0

Final Answer: f(x) = x^2 + 11x + 28